OpenStudy (moongazer):

give the solution set of system 6x + 9y = 7 3x - 6y +14 = 0 Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (saifoo.khan):

-4/3, 5/3

OpenStudy (angela210793):

6x + 9y = 7 6x + 9y = 7 6x+9y=7 3x - 6y +14 = 0 3x - 6y =-14 (*-2)---> -6x+12y=28(add both eq) 21y=35 y=35/21=5/3 replace y in the 1st eq 6x+9*35/21=7 6x=7-15 x=-8/6=-4/3 so the solution is (-4/3;5/3)

OpenStudy (anonymous):

Rearrange each equation to get x (or y) on its own: \[6x + 9y = 7 => 6x = 7 - 9y => x = \frac{7 - 9y}{6}\]\[3x - 6y + 14 = 0 => 3x = 6y - 14 => x = \frac{6y - 14}{3}\] Now we have x in two different forms, which are equal. So set them to be equal! \[\frac{7-9y}{6}=\frac{6y-14}{3}\]\[3(7-9y) = 6(6y-14)\]\[21-27y=36y-84\]\[y=\frac{5}{3}\] Put that y in one of the first equations to find x: \[x=\frac{7-9y}{6}\]\[x=\frac{7-9(5/3)}{6}\]\[x=-\frac{4}{3}\]

OpenStudy (moongazer):

thanks i forgot the negative sign, now i know what is wrong with my solution

OpenStudy (angela210793):

uw

OpenStudy (moongazer):

i'll really be careful know with signs

OpenStudy (moongazer):

my first answer recently is always y= 7

OpenStudy (moongazer):

cause i forgot the - sign

OpenStudy (moongazer):

^_^

OpenStudy (angela210793):

I forget them too :/